### v2

# Learning Diffeomorphism for Image Registration with Time-Continuous Networks using Semigroup Regularization

Abstract
Diffeomorphic image registration (DIR) is a critical task in 3D medical image analysis, aimed at finding topology preserving deformations between pairs of images. Focusing on the solution of the flow map differential equation as the diffeomorphic deformation, recent methods use discrete timesteps along with various regularization terms to penalize the negative determinant of Jacobian and impose smoothness of the solution vector field. In this paper, we propose a novel learning-based approach for diffeomorphic 3D-image registration which finds the diffeomorphisms in the time continuum with only a single regularization term and no additional integration. As one of the fundamental properties of flow maps, we exploit the semigroup property as the only form of regularization, ensuring temporally continuous diffeomorphic flows between pairs of images. Leveraging this property, our method alleviates the need for additional regularization terms and scaling and squaring integration during both training and evaluation. To achieve time-continuous diffeomorphisms, we employ time-embedded UNets, an architecture commonly utilized in diffusion models. The proposed method reveals that ensuring diffeomorphism in a continuous time interval leads to better registration results. Experimental results on four public datasets demonstrate the superiority of our model over both learning-based and optimization-based methods. (https://github.com/mattkia/SGDIR/)
1 Introduction
Diffeomorphic image registration (DIR) is a computational technique used to align images into a common coordinate system using a deformation field, ensuring that the transformation is smooth and invertible [1, 2]. Unlike simpler forms of image registration that might allow folding or tearing, diffeomorphic methods preserve the topological properties of the image since the mapping between corresponding points is continuous, invertible, and differentiable. This property is significant in medical imaging, where anatomical structures need to be aligned accurately without grid folding to facilitate precise analysis and diagnosis.
Addressing the problem of diffeomorphic image registration has led to the development of various approaches, broadly categorized into traditional, optimization-based, and learning-based methods. Traditional methods such as SyN [3], NiftyReg [4], B-Splines [5], and LDDMM [6] usually start with an assumption on the model functional form and try to maximize the similarity between the moving and the fixed images while constraining the model to satisfy some regularity properties. Even though such methods can usually generate fine detailed registrations, they suffer from relatively long computation times. Recent advancements in deep neural networks have yielded optimization-based approaches using neural networks which alleviate the restriction of having an assumption on the model functional and replace the registration model with a neural network with more expressive power [7, 8, 9, 10, 11, 12]. Such models have demonstrated superior performance and lower computation times compared to traditional methods. On the other hand, learning-based approaches [13, 14, 15, 16, 17, 18, 19, 20, 21] have gained traction due to their ability to learn implicit representations of the registration process from large datasets. These methods, still leveraging deep neural network architectures can offer potentially more robust solutions, often significantly reducing computation time compared to traditional and neural optimization-based methods.
Instead of directly modeling the deformation field, most of optimization- and learning-based methods [9, 10, 19] model the velocity field controlling the dynamics of the deformation and use scaling and squaring integration scheme [22, 23] to reconstruct the deformation. Such methods typically regularize their model with various regularization terms to directly penalize the foldings in the deformation and impose smoothness on the vector field gradients and magnitudes. The advent of time-embedded architectures and time-embedded UNets [24, 25, 26] has facilitated the incorporation of the notion of time into the neural networks. Models such as DiffuseMorph [27] directly employ the Diffusion Model architecture to directly and continuously learn the deformation.
In this paper, we present a novel DIR method, called SGDIR (Semi-Group DIR), which is capable of directly generating the continuous deformation without using scaling and squaring integration and multiple regularization terms. Our method is built upon a simple yet important property of flow maps, called the semi-group property [28] which governs the composition rule of flows along time. Using time-embedded UNets our model is able to retrieve the deformation at any time step instantaneously, and hence provide a smooth transition, diffeomorphic at any time step, from the moving image to the fixed image. This is unlike the DiffuseMorph which requires to start from time zero and build up the deformation up to the given time due to following diffusion models paradigm. With a novel regularization term using the semi-group property of flow maps, we ensure diffeomorphic deformations without using any other regularization terms including penalizing negative Jacobian determinants of the deformation. Furthermore, our formulation allows us to simultaneously retrieve the reverse deformation from the fixed image to the moving image.
Our contributions can be summarized as the followings:
  • We introduce a novel DIR method utilizing time-embedded UNets which is capable of producing diffeomophic deformation fields in a time continuum.
  • We remove additional integration schemes such as scaling and squaring integration and regularization terms including penalization of negative Jacobian determinants of the deformation and imposition of smoothness of vector field gradient and magnitude.
  • We introduce semi-group regularization as the only regularization term for leading the model to generate smooth and invertible deformations.
  • We provide theoretical properties and consequences of the semi-group regularization including how it leads to temporally-continuous diffeomorphic deformations.
  • We provide extensive experiments on four 3D image registration datasets including brain and lung MR scans, which demonstrate the outperformance of SGDIR with respect to recent DIR methods.
2 Related Works
2.1 Pairwise Optimization-Based Methods
Traditional methods find the deformation field between a pair of images by minimizing an energy functional while restricting the solution space by imposing some regularity conditions on the model. SyN [6] as a successful traditional model provides a greedy technique for a symmetric diffeomorphic deformation solution by minimizing the cross-correlation between the pair of images. NiftyReg [4] as a powerful package for image registration considers different sets of linear and non-linear (such as B-Splines and Free Form Deformations (FFD)) parametric models [29] and minimizes loss functions such as Normalized Mutual Information (NMI) and Sum of Squared Differences (SSD) to maximize the alignment between the pairs of images. Traditional methods such as Demons and its variants [30, 31, 32] are built upon optical flows. LDDMM [6] is a seminal work in flow-based registration which models the registration as the geodesic paths in the images space and is proper for large deformations while preserving the topology. DARTEL [33] is also another flow-based approach which models the deformations using exponentiated Lie algebras and utilizes the scaling and squaring integration to reconstruct the deformations.
Neural optimization based approaches model the registration using neural networks. IDIR [7] and DNVF [9] are coordinate-based registration methods which utilize implicit neural representations [34]. IDIR directly learns the deformation and applies Jacobian regularization for better topology preservation and hyperelastic [35] regularization for smooth deformation, while DNVF implicitly learns the velocity field and using scaling and squaring integration reconstructs the deformation. DNVF uses Jacobian regularization along with vector field gradient and magnitude regularization. NODEO [8] which also works with voxel coordinate clouds, incorporates neural ODE networks [36] to learn the dynamics of the deformation by estimating the velocity field and finding the deformation using an ODE solver.
2.2 Learning-Based Methods
Unlike optimization-based methods which perform pairwise optimization, learning-based methods try to learn the deformation across large training datasets. The most advantageous aspect of learning-based methods is their short inference time which is orders of magnitude shorter than optimization based methods. As a fundamental work, VoxelMorph [20] directly models the deformation using a UNet and by using localized Normalized Cross Correlation metric and smoothness regularization learns the deformation fields. SYMNet [19] learns the velocity fields of forward and backward deformation and using scaling and squaring integration reconstructs the deformation. SYMNet uses Jacobian and smoothness regularizations along with minimizing the discrepancy between the forward and backward velocity fields, leading to a smoother deformation. LapIRN [37] introduces multi-resolution deformations for better generalization and avoiding local minima, and cLapIRN [16] extends LapIRN and feeds the smoothness regularization factors to the layers of the network. MS-ODE [38] incorporates neural ODEs as a refinement stage by modeling the dynamics of the parameters of a registration model.
DiffuseMorph [27] utilizes denoising diffusion models to learn temporally continuous deformations. It also incorporates Jacobian and smoothness regularization for imposing regularity on the solution. However using diffusion models could be challenging due to expensive training and maintaining visually pleasing outputs. Also, since DiffuseMorph follows the paradigm of diffusion models it requires fixed number of time samples during both training and evaluation. Furthermore, if queried at different time steps, DiffuseMorph requires to build the deformation from time step zero up to the queried time steps. Similar to DiffuseMorph we use time-embedded UNets to learn the temporally continuous deformations, but there are fundamental differences between our method and DiffuseMorph: we use a much simpler UNet without any attention modules; we are training the model with continuous time frame and we can query the model at any time step instantly. Moreover, we only use semigroup regularization introduced in this paper for learning diffeomorphic deformations.
3 Preliminaries
Deformable registration for 3D images seeks for a vector field $\phi:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ which when applied to the moving image $I_{m}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$, deforms it smoothly towards the fixed image $I_{f}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$. $\Omega\subset\mathbb{R}^{3}$ is usually a cubical region. In diffeomorphic image registration we additionally require the deformation field to be an orientation-preserving diffeomorphism, which can be mathematically translated to $\phi$ having positive determinant of Jacobian $|J_{\phi}|$ at all points. The deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ (we will use this definition going forward) can also be considered to be the time-continuous flow map solution of the following autonomous ordinary differential equation [39, 40, 6, 33, 19]:
$\left\{\begin{array}[]{ll}\frac{d\phi}{dt}=\mathbf{v}(\phi)&\\ \phi_{0}(x)=x,&\\ \end{array}\right.$ (3)
where $\phi_{0}(x)$ is short for $\phi(x,t=0)$ and $\mathbf{v}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ is a stationary velocity field governing the dynamics of the flow. In the context of DIR, $\phi_{1}(x)$ is taken to be the deformation which warps $I_{m}$ to $I_{f}$. The utility of the ODE of Eq.(3) is that its solution is guaranteed to be a diffeomorphism. It is shown that considering
$\phi_{\frac{1}{2^{N}}}(x)=x+\frac{\mathbf{v}}{2^{N}}$ (4)
for a relatively large $N$ (e.g., $N=8$) allows one to use the scaling and squaring integration scheme to find $\phi_{1}(x)$ by iteratively applying the deformations [22, 23]:
$\phi_{1/2^{N-1}}=\phi_{1/2^{N}}\circ\phi_{1/2^{N}}\Rightarrow\phi_{1/2^{N-2}}=\phi_{1/2^{N-1}}\circ\phi_{1/2^{N-1}}\Rightarrow\dots\Rightarrow\phi_{1}=\phi_{1/2}\circ\phi_{1/2}.$ (5)
Conventional DIR methods, taking advantage of the above properties, assume the following general loss function to address the diffeomorphic registration problem
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D}}\left[\mathcal{L}_{\text{sim}}(I_{f},\phi[I_{m}])+\mathcal{L}_{\text{reg}}(\phi)\right],$ (6)
where $\mathcal{D}$ is the distribution of the images, $\mathcal{L}_{\text{sim}}(.,.)$ is a measure of similarity between the fixed image and the warped moving image, and $\mathcal{L}_{\text{reg}}(\phi)$ contains restrictive regularization terms over the deformation $\phi$. Here, $\phi[I]$ means the warping of image $I$ under the deformation field $\phi$. Starting with an estimation of the vector field $\mathbf{v}$ in Eq.(4), the deformation $\phi$ (more precisely, $\phi_{1}$) is obtained by scaling and squaring integration, and the whole model is optimized to minimize the loss Eq. (6).
On the other hand, a necessary and sufficient condition of $\phi$ as the flow map solution of the ODE of Eq.(3) is that it satisfies the semigroup property; i.e., for any time steps $t$ and $s$ it holds [28]:
$\phi(x,0)=x~{}\text{and}~{}\phi(\phi(x,s),t)=\phi(\phi(x,t),s)=\phi(x,s+t),$ (7)
or in short
$\phi_{0}=Id~{}\text{and}~{}\phi_{t}\circ\phi_{s}=\phi_{s}\circ\phi_{t}=\phi_{t+s},$ (8)
where $Id$ denotes identity function. Intuitively, semigroup property implies that deforming an image $I$ up to a time $t+s$ is equivalent to first deforming the image up to time $t$ to obtain $I_{t}$, and then applying the deformation again up to time $s$ given that the initial condition has changed to $I_{t}$. Also, since $\phi_{t}\circ\phi_{-t}(x)=\phi_{-t}(x)\circ\phi_{t}=\phi_{0}(x)=x$, we deduce that
$\phi^{-1}_{t}(x)=\phi_{-t}(x),$ (9)
for all values of $x$. Given that $\phi_{t}$ is a differentiable function, Eq. (9) guaranties the bijectivity of $\phi_{t}(x)$, and hence, the diffeomorphic property of $\phi_{t}$. In the next section we’ll show how we exploit the semigroup property of Eq.(8) to circumvent the scaling and squaring integration and other regularization terms and perform the registration in a continuous time frame in contrast to former DIR methods which find the deformation at time $1$.
4 Proposed Method
Here we elaborate our method, dubbed as SGDIR. Let $I_{f},I_{m}:\Omega\longrightarrow\mathbb{R}$ be the fixed and moving images, respectively, sampled from the images distribution $\mathcal{D}$. $\Omega\subset\mathbb{R}^{3}$ is a $D\times H\times W$ cubical region. We are looking for a time-continuous deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ to deform the moving image towards the fixed image, and vice versa. We assume that $\phi$ is the flow map solution of the ODE (3), for some unknown stationary vector field $\mathbf{v}$, and therefore, it must satisfy the semigroup property of Eq.(8). If we warp $I_{m}$ up to time $t$ using the deformation $\phi_{t}$, and warp $I_{f}$ up to time $1-t$ using the inverse deformation $\phi^{-1}_{1-t}$, we must reach to a same point due to the continuity of the trajectory of $\phi$. Based on Eq.(9) on the inversion of $\phi$, we can write
Figure 1: The general schematic of the SGDIR training. As illustrated in the figure, there are no additional integration or extra loss functions applied which makes the whole procedure simple. During inference, we can query the model for any desired $t$, e.g., $t=1$ for the final deformation.
$\phi^{-1}_{1-t}[I_{f}]=\phi_{t-1}[I_{f}]=\phi_{t}[I_{m}].$ (10)
Eq.(10) allows us to define time-continuous similarity loss. For measuring the similarity, we use the localized Normalized Cross Correlation (NCC) between the images [20, 8] (see Appendix 7.1).
In this work we use a local window of size $11$ in NCC. Using Eq.(10), we introduce a time-continuous similarity loss:
$\mathcal{L}^{t}_{\text{sim}}$ $=-NCC(\phi_{t-1}[I_{f}],\phi_{t}[I_{m}]),$ $~{}\forall t\in[0,1].$ (11)
We propose to learn the time-continuous deformation $\phi(x,t)$ using a time-embedded UNet as shown in Figure 1. The Time-Embedded UNet (TE-UNet), frequently used in diffusion models, is capable of incorporating the notion of time and is suitable for learning flow maps. To achieve this goal, we model the deformation $\phi$ as
$\phi_{t}(x;\theta)=\phi(x,t;\theta)=x+t\mathbf{F}(x,t;I_{f},I_{m},\theta),$ (12)
for all $t\in[-1,1]$, where $\mathbf{F}(x,t;I_{f},I_{m},\theta)$ is a TE-UNet with learnable parameters $\theta$ which also receives a pair of fixed and moving images. The reason behind the specific formulation of Eq.(12) is that we can make sure that at $t=0$ we have $\phi_{0}(x)=x$, hence satisfying the identity initial condition of the ODE (3). In order for $\phi$ to be a valid flow map, it needs to satisfy the semigroup property of Eq. (8). In practice, forcing the model to perfectly satisfy the semigroup property could be cumbersome and expensive, since we need to sample $t$ and $s$ independently and impose Eq. (11) twice, which requires more forward calls from the network. To tackle this issue we propose the following regularization,
$\mathcal{L}^{t}_{\text{reg}}=\|\phi_{2t-1}-\phi_{t}\circ\phi_{t-1}\|_{2}+\|\phi_{2t-1}-\phi_{t-1}\circ\phi_{t}\|_{2},$ (13)
for all $t\in[0,1]$, and we show that it suffices to render $\phi$ invertible and consequently a diffeomorphism at all time steps $t\in[-1,1]$. In Eq. (13), .$\|.\|_{2}$ denotes the $L_{2}$-norm, and forces the model to satisfy the semigroup property for $t$ and $s=t-1$. The second term in Eq. (13) is added only to impose symmetry and provide diffeomorphism in $[-1,0]$ interval. Similar to all DIR methods we implement the composition operation by trilinear interpolation of grids over each other [19, 9, 37, 16]. From Eq. (13), one could instantly say that a model that minimizes Eq. (13), will satisfy the following rules
$\phi_{2t-1}=\phi_{t}\circ\phi_{t-1},~{}\phi_{2t-1}=\phi_{t-1}\circ\phi_{t}.$ (14)
Proposition 1.
The deformation $\phi(x,t)$ optimally learned by minimizing Eq. (11) and Eq. (13), is a diffeomorphism at any time $t\in[-1,1]$.
Proof.
See Appendix 7.2. ∎
Finally, we can propose the following loss function which is in concordance with Eq. (6):
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D},t\sim Uni(0,1)}\left[\mathcal{L}^{t}_{\text{sim}}+\lambda\mathcal{L}^{t}_{\text{reg}}\right],$ (15)
where $Uni(0,1)$ is the uniform distribution on $[0,1]$ and $\mathcal{L}^{t}_{\text{sim}}$ and $\mathcal{L}^{t}_{\text{reg}}$ are defined as in Eq. (11) and Eq. (13), respectively. $\lambda$ is the regularization factor controlling the level of diffeomorphism. In the training phase, we randomly sample a pair of images to be registered along with a time step uniformly taken from the interval $[0,1]$ to minimize Eq. (15). The trained model is capable of warping either $I_{f}$ or $I_{m}$ towards each other at any time step. Thus, the model is more versatile compared to the most previous works that provide the deformation in only one direction.
It is also worth mentioning that some prior works [41, 42] enforce forward-backward deformation consistency as a regularization term by constructing a time-independent forward deformation only at the trajectory endpoint ($\phi_{1}$), while explicitly modeling the backward deformation ($\phi_{1}^{-1}$). In contrast, SGDIR allows evaluating the deformation field $\phi_{t}$ at arbitrary time steps, without separately modeling the inverse. Consistency between forward and backward deformations, as we prove, is obtained as a natural result of semigroup regularization, rather than through direct imposition. LDDMM [6] models a time-dependent deformation using discrete time steps as opposed to our continuous time steps and uses forward-backward calculations to compute the velocity field explicitly. In contrast, we do not compute any velocity field explicitly.
5 Experiments
5.1 Datasets and Preprocessing
To evaluate SGDIR and compare its performance with other methods, we consider four publicly available datasets on 3D MR brain and lung scans widely used in the literature, namely OASIS [43], CANDI [44], LPBA40 [45], and DIRLAB [46]. OASIS dataset contains 416 T1 weighted scans from subjects aging from 18 to 96 with 100 of them diagnosed with mild to moderate Alzheimer’s disease. The segmentation masks of 35 subcortical regions available in the dataset serve as the ground truth for further evaluation of the registration. CANDI dataset contains T1 weighted brain scans of 117 subjects divided into 4 different subgroups including Healthy Control (HC), Schizophrenia Spectrum (SS), Bipolar Disorder with Psychosis (BPDwithPsy), and Bipolar Disorder without Psychosis (BPDwithoutPsy). The segmentation masks of 32 subcortical regions available in the dataset are used as the ground truth for the evaluation. LPBA40 dataset comes with 40 scans and the segmentation maps of 56 cortical and subcortical regions. For these datasets we use the skull stripped, MNI152 1mm normalized images and center-crop the images to the size $160\times 192\times 144$. For all datasets we use all the available anatomical structures for evaluation. DIRLAB dataset contains 10 pairs of lung CT scans corresponding to 10 patients. Each pair consists of the scans of lungs in the extreme inhaling and exhaling phases. The dataset also comes with 300 landmark points per image, annotated manually by experts, which are used for the evaluation.
We use min-max intensity normalization to transform the voxel intensities into the $[0,1]$ range. In the training phase, in order for the arguments of Proposition 1 remain valid we add a slight standard Gaussian noise with 0 mean and 0.01 standard deviation to the data. This makes sure that no pairs of voxels have precisely the same intensity levels while the image pair remaining visually intact.
5.2 Experimental Settings
In the OASIS dataset we use the first 256 subjects for training the model, the next 50 subjects for the validation and the rest of the subjects for the test set. By random sampling we generate 900 pairs for training, 100 pairs for validation and 1000 pairs for test. In the CANDI dataset, we mix all the subgroups and then use 80 subjects for training, 11 subjects for validation, and the rest for the test set. In a similar way to the OASIS dataset, we generate 400 pairs for training, 25 pairs for validation, and 300 pairs for test. For the LPBA40 dataset 190 pairs from 20 subjects are used as the training set, 10 pairs from 5 subjects are used as the validation set and 105 pairs from 15 subjects are used for the test set. For the DIRLAB dataset, 8 subjects are used for training and 2 subjects are used for test. We compare the performance of our model in terms of Dice score (for the OASIS, CANDI, and LPBA40 datasets), TRE (for the DIRLAB dataset), and the percentage of negative determinants of Jacobian of the deformation with a traditional optimization-based method, SyN [3], as one of the most successful methods for diffeomorphic image registration, learning-based methods including SYMNet [19], VoxelMorph [20], LapIRN [37], and DiffuseMorph [27], and optimization-based methods such as IDIR [7], NODEO [8], DNVF and its variant Cas-DNVF [9]. For SyN, we used the implementation in the DIPY package [47] and used cross-correlation for the loss function following the convention of the VoxelMorph. For other methods, except for DNVF/Cas-DNVF whose implementation was not available at the time of writing this paper, we used the codes available in their corresponding official repositories. The architecture of the UNet along with the time embedding modules is shown in Figure 1. The UNet encoder has 4 down-sampling layers of dimensions 32, 64, 128, 128, and 256, and the decoder has 4 up-sampling layers with the same dimensions as the down-sampling layers but in the reversed order. The time-embedding dimension is set to 64. All the activation functions for the layers are set to SiLU [48] to provide more smoothness to the network. In our experiments the regularization factor for the semigroup term is set to $\lambda=5e^{5}$, the NCC window size is set to 11, and the training is carried out for 100 epochs using Adam optimizer with the learning rate $lr=1e^{-4}$. The whole implementation is done in PyTorch and is tested on both NVIDIA GTX 1080 Ti and NVIDIA Tesla V100 GPUs with 32Gb RAM.
5.3 Evaluation Metrics
Following the convention of evaluating diffemorphic image registration methods ([9, 8, 19, 37]), we use Dice Similarity Coefficient to measure the similarity of the registered and target segmentation maps for the brain scan dataset, and we use TRE metric to measure the error between the registered and target landmark points for the DIRLAb dataset. We also use the negative Jacobian determinant percentage to measure the amount of disturbance in the topology of the moving image.
To compute the Dice score, we follow the convention of VoxelMorph [20] for computing the Dice score, which measures the mean Dice over all anatomical structures (35, 32, and 56 anotimical structures for the OASIS, CANDI, and LPBA40 datasets, respectively). For the DIRLAB dataset, the images are resampled to isotropic voxels of size $1mm$, and the TREs are reported in terms of average error in millimeter.
5.4 Results
Tables 1 and 2 provide a comprehensive comparison of the performance of SGDIR with other aforementioned methods on OASIS, CANDI, LPBA40, and DIRLAB datasets, respectively. Table 1 provides the performance of models in terms of Dice score and the percentage of negative Jacobian determinants. All the values are expressed as their average over the test set along with their standard deviation. For DNVF and Cas-DNVF due to unavailability of the implementation the results are directly taken from the corresponding paper. We consider the percentage of negative Jacobian determinants smaller than $10^{-5}\%$ negligible and show it as $\approx 0$ in the tables. Also, Figure 2 illustrates the qualitative comparison between our model and several other models on the OASIS dataset in terms of the warped image, warped segmentations, and the warping grids. Figure 3 illustrates the visual comparison between SGDIR and IDIR over the DIRLAB dataset.
Table 1 indicates that SGDIR outperforms competent methods both in terms of Dice score, and $|det(J_{\phi})<0|$, and Table 2 demonstrates the advantage of SGDIR in terms of the TRE metric. This implies that the imposition of the semigroup property as a regularization in the time continuum of the deformation field can in fact lead to a smooth deformation.
Among the compared methods, DiffuseMorph is also capable of generating a temporal deformation flow. Therefore, we have provided a visual comparison of warped images within 7 time steps for the OASIS dataset in Figure 4. In this example we have used the subject id 1 and 2 of the OASIS dataset as the fixed image and the moving image respectively. More visual results are also provided in Appendix 7.3.
Table 1: Comparison of the performance of SGDIR with other methods on the OASIS, CANDI, and LPBA40 datasets. The values inside the parenthesis indicate the standard deviation. Cells with - indicate that the corresponding information is missing in the literature. $\approx 0$ sign is used for negative determinant percentages smaller than an order of $10^{-5}$.
OASIS CANDI LPBA40
Category Model Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$) Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$) Dice ($\uparrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 78.14 (1.32) 0.0531% (0.0241) 73.60 (1.15) 0.0011% (0.0008) 66.82 (3.17) 0.0194% (0.0072)
Learning VoxelMorph 75.93 (0.044) 0.0026% (0.0027) 75.63 (1.78) 0.0011% (0.0010) 70.72 (2.33) 0.0023% (0.0014)
SYMNet 79.39 (1.74) 0.0043% (0.0017) 75.89 (1.13) 0.0015% (0.0013) 69.98 (1.37) 0.0018% (0.0012)
LapIRN 78.80 (1.92) 0.0011% (0.0003) 76.23 (1.76) 0.0004% (0.0003) 71.88 (1.36) $\mathbf{\approx 0}$
DiffuseMorph 76.23 (2.23) 0.0551% (0.0431) 75.63 (2.01) 0.0041% (0.0012) 68.73 (2.88) 0.0032% (0.0011)
Optimization IDIR 72.31 (1.239) 0.0389% (0.0025) 74.42 (1.12) 0.0113% (0.0081) 67.93 (2.03) 0.0013% (0.0008)
NODEO 80.86 (1.14) 0.0105% (0.0115) 78.02 (1.52) $\approx$ 0% 72.31 (1.12) 0.0009% (0.0004)
DNVF 79.4 0.0015% 78.6 0.0003% - -
Cas-DNVF 81.5 0.0036% - - - -
Our Method SGDIR 82.23 (1.3851) 0.0009% (0.0007) 80.73 (1.25) $\mathbf{\approx 0}$74.23 (2.15) 0.0007% (0.0003)
Figure 2: Qualitative comparison of SGDIR with other methods on the OASIS dataset in terms of the warped images (first row), warped segmentations (second row) and warping grids (third row).
Table 2: Comparison of the performance of SGDIR with other methods on the DIRLAB dataset. TRE is reported in terms of millimeters which is the same as the voxel size
Category Model TRE (mm) ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 1.35 0.0034%
Learning VoxelMorph 1.33 0.0011%
SYMNet 1.27 0.0006%
LapIRN 1.21 0.0001%
DiffuseMorph 1.35 0.0010%
Optimization IDIR 1.13 0.0005%
NODEO 1.08 0.0003%
Our Method SGDIR 0.98 0.0003%
Figure 3: Visual comparison of SGDIR with IDIR over the DIRLAB dataset. Deformation fields are depicted over the landmark points inside the lungs.
Figure 4: The visual comparison of the evolution of the deformation between DiffuseMorph and SGDIR on the OASIS dataset over 7 time steps.
5.5 Ablation Study
In addition to the main results on the efficacy of SGDIR in providing high Dice scores while preserving the topology, we investigate the effect of some hyperparameters and design choices along with some consequent properties of SGDIR. More precisely, we investigate the effect of the semigroup regularization factor $\lambda$, the effect of continuous time scale against discrete time, and the topology-preserving property throughout the time interval. Since the behaviour of the model is the same for all datasets, we provide the analysis only for the OASIS dataset which is a more complex dataset with larger deformations.
5.5.1 Effect of Regularization Factor
Eq. (13) is responsible for imposing the semigroup property to the model and consequently, make the model produce topology-preserving non-folding deformations. Table 3 provides the effect of changing the factor $\lambda$ for the semigroup regularization on the Dice score and $|det(J_{\phi})<0|$. This table indicates that for a choice of a very large $\lambda$ we could achieve zero-folding deformations with a decrease in the dice score. Decreasing $\lambda$ gives more freedom to the model to increase the similarity of the deformed image and the fixed image in exchange for a larger folding percentage. An interesting property exhibited here is that drastically decreasing $\lambda$ (e.g., $\lambda=1$) is not helpful for achieving the highest Dice score. This indicates that the semigroup property is also helpful in narrowing down the trajectory space of the deformations, helping the model to converge to the more optimal solution.
Table 3: Analysis of the effect of the semigroup regularization factor on the Dice score and the percentage of negative Jacobian determinants over the OASIS dataset.
$\lambda$ Dice ()($\uparrow$)( ↑ ) $|\text{det}(J_{\phi})<0|$ ()($\downarrow$)( ↓ )
$10^{7}$ 76.20 0.0000%
$10^{5}$ 82.11 0.0011%
$10^{4}$ 83.45 3.0691%
$10^{3}$ 82.66 6.0955%
$10^{2}$ 80.01 6.7403%
$10$ 80.23 7.0794%
$1$ 79.80 7.8612%
Figure 5: The percentage of negative Jacobian determinants through time. The upper figure compares the effect of $\lambda$ on the $|det(J_{\phi})<0|$ and the bottom s figure focuses on the best configuration model. The shaded area in the bottom figure indicates the variance across the test set.
5.5.2 Diffeomorphism Through Time
Applying the semigroup regularization ensures that the deformation is a diffeomorphism not only at the endpoints, but also at all times $t\in[0,1]$. Figure 5 illustrates how $|det(J_{\phi})<0|$ varies across time. The negative time interval is used for reverse warping; i.e., warping the fixed image towards the moving image. The upper figure illustrates the effect of different values of $\lambda$ on $|det(J_{\phi})<0|$ throughout the time interval, and one can see that a proper factor for the semigroup regularization could lead to an almost perfect diffeomorphism in each time step. The bottom figure, focuses on the performance of the best model. The purple plot and the right y-axis indicate the Dice score of intermediate warped images throughout time and the black plot and the left y-axis indicate the $|det(J_{\phi})<0|$. The shaded areas for both plots show the variance of the corresponding metric. We observe that the semigroup regularization successfully preserves the topology at all time steps, while keeping the Dice score high during the transition.
Table 4: Analysis of the effect of using discrete time steps on the Dice score and the percentage of negative Jacobian determinants.
# Time Steps Dice $|\text{det}(J_{\phi})<0|$
00 73.12 13.121%
$1$ 73.68 0.152%
$2$ 74.23 0.065%
$4$ 75.03 0.008%
$8$ 75.45 0.002%
$\infty$ 82.23 0.0009%
5.5.3 Effect of Discrete Time Steps
SGDIR can compute the deformation at an arbitrary time step using a continuous sampling from the time interval $[0,1]$. Here, we investigate the effect of using discrete time steps for training the model. The number of time steps indicates the number of intermediate time steps used during training. For instance, 0 time steps means that only the endpoint are used for registration (at time 1), and 8 time steps means that 8 intermediate time steps is used during training. $\infty$ is the continuous time sampling as done in our method. At the evaluation phase, we query the model at times 1 and -1 to warp the moving image towards the fixed image, and vice versa. Table 4 demonstrates that increasing the number of time samples in the interval $[0,1]$ improves the performance of the model both in terms of Dice score and $|det(J_{\phi})<0|$, while the best performance is achieved by continuous sampling. Figure 6 in Appendix 7.3 provides an illustrative inspection of this effect as well.
6 Conclusion
In this paper, we proposed SGDIR, a novel learning-based model for diffeomorphic image registration based on time-embedded UNet architecture and using semigroup-based regularization. The time-embedded network allows us to compute deformation and consequently warp images at any time point. The semigroup regularization enables the model to generate diffeomorphic deformations and its inverse. The resulting model is capable of generating smooth warpings from the moving image to the fixed image and vice versa without requiring any integration scheme or using further regularization terms on the smoothness of the deformation. We evaluated our model on four benchmark datasets on 3D medical image registration and demonstrated superiority in various metrics compared to the state-of-the-art optimization and learning-based methods. In a broader perspective, the proposed framework provides a systematic method for addressing problems which can be modeled as an autonomous ODE.