### v3

# Learning Diffeomorphism for Image Registration with Time-Continuous Networks using Semigroup Regularization

Abstract
Diffeomorphic image registration (DIR) is a fundamental task in 3D medical image analysis that seeks topology-preserving deformations between image pairs. To ensure diffeomorphism, a common approach is to model the deformation field as the flow map solution of a differential equation, which is solved using efficient schemes such as scaling and squaring along with multiple smoothness regularization terms. In this paper, we propose a novel learning-based approach for diffeomorphic 3D image registration that models diffeomorphisms in a continuous-time framework using only a single regularization term, without requiring additional integration. We exploit the semigroup property—a fundamental characteristic of flow maps—as the sole form of regularization, ensuring temporally continuous diffeomorphic flows between image pairs. Leveraging this property, we prove that our formulation directly learns the flow map solution of an ODE, ensuring continuous inverse and cycle consistencies without explicit enforcement, while eliminating additional integration schemes and regularization terms. To achieve time-continuous diffeomorphisms, we employ time-embedded UNets, an architecture commonly used in diffusion models. Our results demonstrate that modeling diffeomorphism continuously in time improves registration performance. Experimental results on four public datasets demonstrate the superiority of our model over state-of-the-art diffeomorphic methods. Additionally, comparison to several recent non-diffeomorphic deformable image registration methods shows that our method achieves competitive Dice scores while significantly improving topology preservation. (https://github.com/mattkia/SGDIR/)
1 Introduction
Diffeomorphic image registration (DIR) is a computational technique used to align images into a common coordinate system using a deformation field, ensuring that the transformation is smooth and invertible [1, 2, 3]. Unlike simpler forms of image registration that might allow folding or tearing, diffeomorphic methods preserve the topological properties of the image since the mapping between corresponding points is continuous, invertible, and differentiable. This property is crucial in medical imaging, where anatomical structures need to be aligned accurately without grid folding to facilitate precise analysis and diagnosis.
Addressing the problem of diffeomorphic image registration has led to the development of various approaches, broadly categorized into traditional, instance optimization-based, and learning-based methods. Traditional methods such as SyN [4], NiftyReg [5], B-Splines [6], and LDDMM [7] usually start with an assumption on the model functional form and try to maximize the similarity between the moving and the fixed images while constraining the model to satisfy some regularity properties. Even though such methods can usually generate fine-detailed registrations, they suffer from relatively long computation times. Recent advances in deep neural networks have yielded instance optimization-based approaches using neural networks which alleviate the restriction of having an assumption on the model functional and replace the registration model with a neural network with more expressive power [8, 9, 10, 11, 12, 13]. Such models have demonstrated superior performance and lower computation times compared to traditional methods. On the other hand, learning-based approaches [14, 15, 16, 17, 18, 19, 20, 21] have gained traction due to their ability to learn implicit representations of the registration process from large datasets. These methods, still leveraging deep neural network architectures can offer potentially more robust solutions, often significantly reducing inference time compared to traditional and neural optimization-based methods.
Instead of directly modeling the deformation field, most of optimization- and learning-based methods [10, 11, 20] model the velocity field controlling the dynamics of the deformation and use scaling and squaring integration scheme to reconstruct the deformation [22, 23]. Such methods typically regularize their model with various regularization terms to directly penalize the foldings in the deformation and impose smoothness on the vector field gradients and magnitudes. The advent of time-embedded architectures and time-embedded UNets [24, 25, 26] has facilitated the incorporation of the notion of time into the neural networks. Models such as DiffuseMorph and DiffuseReg [27, 28] directly employ the Diffusion Model architecture to directly and continuously learn the deformation.
It is also known that the inverse consistency of a registration model plays a critical point in the behavior of the registration method especially for the medical image registration task [29]. To this end, various methods have incorporated cycle and inverse consistencies either into the model’s architecture or in their training procedure [29, 30, 31, 32].
In this paper, we present a novel DIR method, called SGDIR (Semi-Group DIR), which is capable of directly generating the continuous deformation without using scaling and squaring integration and multiple regularization terms. Our method is built upon a simple yet important property of flow maps, called the semigroup property [33] which governs the composition rule of flows along time. Using time-embedded UNets our model is able to retrieve the deformation at any time step instantaneously, and hence provides a smooth transition, diffeomorphic at any time step, from the moving image to the fixed image, and vice versa. We will prove that our particular semigroup regularization ensures that the model learns a deformation field that is indeed the solution to an ODE, hence, implicitly satisfying the diffeomrophic property and cycle and inverse consistencies, without using any other regularization terms including penalizing negative Jacobian determinants of the deformation. Our contributions can be summarized as the followings:
  • We introduce a novel DIR method utilizing time-embedded UNets which is capable of producing diffeomophic deformation fields in a time continuum.
  • We introduce a simple semigroup regularization as the only regularization term and we mathematically prove that our regularization term leads the model towards learning the solution of the deformation ODE, hence, implicitly satisfying the diffeomorphic property.
  • We remove additional integration schemes such as scaling and squaring integration and regularization terms including penalization of negative Jacobian determinants of the deformation and imposition of smoothness of vector field gradient and magnitude.
  • We provide extensive experiments on four 3D image registration datasets on brain MR scans, which demonstrate the superior accuracy of SGDIR compared to recent DIR methods. When compared with recent non-diffeomorphic deformable methods, SGDIR show significant improvement in topology preservation without sacrificing accuracy.
2 Related Works
2.1 Pairwise Optimization-Based Methods
Traditional methods find the deformation field between a pair of images by minimizing an energy functional while restricting the solution space by imposing some regularity conditions on the model. SyN [7] as a successful traditional model provides a greedy technique for a symmetric diffeomorphic deformation solution by minimizing the cross-correlation between the pair of images. NiftyReg [5] as a powerful package for image registration considers different sets of linear and non-linear parametric models [34] and minimizes loss functions such as Normalized Mutual Information (NMI) and Sum of Squared Differences (SSD) to maximize the alignment between the pairs of images. Traditional methods such as Demons and its variants [35, 36, 37] are built upon optical flows. LDDMM [7] is a seminal work in flow-based registration that models the registration as the geodesic paths in the images space and is proper for large deformations while preserving the topology. DARTEL [38] is also another flow-based approach which models the deformations using exponentiated Lie algebras and utilizes scaling and squaring integration to reconstruct the deformations.
Neural optimization based approaches model registration using neural networks. IDIR [8], DNVF [10], and NePhi [39] are coordinate-based registration methods which utilize implicit neural representations [40]. IDIR directly learns the deformation and applies Jacobian regularization for better topology preservation and hyperelastic regularization for smooth deformation [41], while DNVF implicitly learns the velocity field and using scaling and squaring integration, reconstructs the deformation. DNVF uses Jacobian regularization along with vector field gradient and magnitude regularization. NODEO [9] which also works with voxel coordinates, incorporates neural ODEs [42] to learn the dynamics of the deformation by estimating the velocity field and finding deformation using an ODE solver.
2.2 Learning-Based Methods
Unlike optimization-based methods which perform pairwise optimization, learning-based methods try to learn the deformation across large training datasets. The most advantageous aspect of learning-based methods is their short inference time which is orders of magnitude shorter than that of optimization based methods. As a pioneering work, VoxelMorph [21] directly models the deformation using a UNet and by using localized Normalized Cross Correlation metric and smoothness regularization learns the deformation fields. SYMNet [20] learns the velocity fields of forward and backward deformation and using scaling and squaring integration reconstructs the deformation. SYMNet uses Jacobian and smoothness regularizations along with minimizing the discrepancy between the forward and backward velocity fields, leading to a smoother deformation. LapIRN [43] and PULPo [44] employ multi-resolution deformations for better generalization and avoiding local minima. cLapIRN [17] extends LapIRN and feeds the smoothness regularization factors to the layers of the network. MS-ODE [45] incorporates neural ODEs as a refinement stage by modeling the dynamics of the parameters of a registration model. TransMorph [46] introduces a general deformable image registration method by utilizing Vision Transformers, and further they customize their model for diffeomorphic registration, known as TransMorph-diff. H-ViT also uses the vision transformers along with cross-attention mechanism to produce more accurate deformations [47].
DiffuseMorph [27] utilizes denoising diffusion models to learn temporally continuous deformations. It also incorporates Jacobian and smoothness regularizations for imposing regularity on the solution. Instead of focusing on the images, DiffuseReg denoises the deformation field and follows the diffusion models strategy [28]. However, using diffusion models could be challenging due to expensive training, maintaining visually pleasing outputs, and slow generation. We also use time-embedded UNets to learn the temporally continuous deformations. Unlike the diffusion model paradigm, we are able to output the deformation instantly at any timestep instead of building it from time step zero. Our simple design endows us with the ability to generate the deformation fields significantly faster.
3 Preliminaries
Figure 1: Schematic of SGDIR. As illustrated in the figure, there are no additional integration or extra loss functions to make the whole training procedure simple. During inference, we can query the model for any desired $t$, e.g., $t=1$ for the final deformation.
Deformable registration for 3D images seeks for a vector field $\phi:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ that when applied to the moving image $I_{m}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$, deforms it smoothly towards the fixed image $I_{f}:\Omega\subset\mathbb{R}^{3}\longrightarrow\mathbb{R}$. The domain $\Omega\subset\mathbb{R}^{3}$ is usually a cubical region. In diffeomorphic image registration, we additionally require the deformation field to be an orientation-preserving diffeomorphism, which can be mathematically translated to $\phi$ having positive determinant of Jacobian $|J_{\phi}|$ at all points. The deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ (we will use this definition going forward) can also be considered to be the time-continuous flow map solution of the following autonomous ordinary differential equation [48, 49, 7, 38, 20]:
$\left\{\begin{array}[]{ll}\frac{d\phi}{dt}=\mathbf{v}(\phi)&\\ \phi_{0}(x)=x,&\\ \end{array}\right.$ (3)
where $\phi_{0}(x)$ is short for $\phi(x,t=0)$ and $\mathbf{v}:\mathbb{R}^{3}\longrightarrow\mathbb{R}^{3}$ is a stationary velocity field that governs the dynamics of the flow. In the context of DIR, $\phi_{1}(x)$ is taken to be the deformation that warps $I_{m}$ to $I_{f}$. The utility of the ODE (3) is that its solution is guaranteed to be a diffeomorphism. It is shown that considering
$\phi_{\frac{1}{2^{N}}}(x)=x+\frac{\mathbf{v}}{2^{N}}$ (4)
for a relatively large $N$ (e.g., $N=8$) allows one to use the scaling and squaring integration scheme to find $\phi_{1}(x)$ by iteratively applying the deformations [22, 23]:
$\phi_{\frac{1}{2^{N-1}}}=\phi_{\frac{1}{2^{N}}}\circ\phi_{\frac{1}{2^{N}}}\Rightarrow\phi_{\frac{1}{2^{N-2}}}=\phi_{\frac{1}{2^{N-1}}}\circ\phi_{\frac{1}{2^{N-1}}}\Rightarrow\dots\Rightarrow\phi_{1}=\phi_{1/2}\circ\phi_{1/2}.$ (5)
Conventional DIR methods, taking advantage of the above properties, assume the following general loss function to address the diffeomorphic registration problem.
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D}}\left[\mathcal{L}_{\text{sim}}(I_{f},\phi[I_{m}])+\mathcal{L}_{\text{reg}}(\phi)\right],$ (6)
where $\mathcal{D}$ is the distribution of the training images, $\mathcal{L}_{\text{sim}}(.,.)$ is a measure of similarity between the fixed image and the warped moving image, and $\mathcal{L}_{\text{reg}}(\phi)$ contains restrictive regularization terms over the deformation $\phi$. Here, $\phi[I]$ means the warping of the image $I$ under the deformation field $\phi$. Starting with an estimation of the vector field $\mathbf{v}$ in Eq. 4, the deformation $\phi$ (more precisely, $\phi_{1}$) is obtained by scaling and squaring integration, and the whole model is optimized to minimize the loss Eq. 6.
On the other hand, a necessary and sufficient condition of $\phi$ as the flow map solution of the ODE (3) is that it satisfies the semigroup property; i.e., for any time steps $t$ and $s$ it holds [33]:
$\phi(x,0)=x~{}\text{and}~{}\phi(\phi(x,s),t)=\phi(\phi(x,t),s)=\phi(x,s+t),$ (7)
or in short,
$\phi_{0}=Id~{}\text{and}~{}\phi_{t}\circ\phi_{s}=\phi_{s}\circ\phi_{t}=\phi_{t+s},$ (8)
where $Id$ denotes identity function. Intuitively, semigroup property implies that deforming an image $I$ up to a time $t+s$ is equivalent to first deforming the image up to time $t$ to obtain $I_{t}$, and then applying the deformation again up to time $s$ given that the initial condition has changed to $I_{t}$. Also, since $\phi_{t}\circ\phi_{-t}(x)=\phi_{-t}\circ\phi_{t}(x)=\phi_{0}(x)=x$, we deduce that
$\phi^{-1}_{t}(x)=\phi_{-t}(x),$ (9)
for all values of $x$. Given that $\phi_{t}$ is a differentiable function, Eq. 9 guaranties the bijectivity of $\phi_{t}(x)$, and hence, the diffeomorphic property of $\phi_{t}$. In the next section we’ll show how we exploit the semigroup property of Eq. 8 to circumvent the scaling and squaring integration and other regularization terms and perform the registration in a continuous time frame in contrast to former DIR methods which find the deformation at time $1$.
4 Proposed Method
Here we elaborate our method, dubbed as SGDIR. Let $I_{f},I_{m}:\Omega\longrightarrow\mathbb{R}$ be the fixed and moving images, respectively, sampled from the image distribution $\mathcal{D}$. The domain $\Omega\subset\mathbb{R}^{3}$ is a $D\times H\times W$ cubical region. We are looking for a time-continuous deformation field $\phi:\Omega\times\mathbb{R}\longrightarrow\mathbb{R}^{3}$ to deform the moving image towards the fixed image, and vice versa. We assume that $\phi$ is the flow map solution of the ODE (3), for some unknown stationary vector field $\mathbf{v}$, and therefore, it must satisfy the semigroup property of Eq. 8. If we warp $I_{m}$ up to time $t$ using the deformation $\phi_{t}$, and warp $I_{f}$ up to time $1-t$ using the inverse deformation $\phi^{-1}_{1-t}$, we must reach to a same point due to the continuity of the trajectory of $\phi$. Based on Eq. 9, we can write
$\phi^{-1}_{1-t}[I_{f}]=\phi_{t-1}[I_{f}]=\phi_{t}[I_{m}].$ (10)
Eq. 10 allows us to define a time-continuous similarity loss. For measuring the similarity, we use the localized Normalized Cross Correlation (NCC) between the images [21, 9] (see Appendix 7.3). In this work we use a local window of size $11$ for NCC. Using Eq. 10, we introduce a time-continuous similarity loss:
$\mathcal{L}^{t}_{\text{sim}}$ $=-NCC(\phi_{t-1}[I_{f}],\phi_{t}[I_{m}]),$ $~{}\forall t\in[0,1].$ (11)
We propose to learn the time-continuous deformation $\phi(x,t)$ using a time-embedded UNet as shown in Figure 1. The Time-Embedded UNet (TE-UNet), frequently used in diffusion models, is capable of incorporating the notion of time and is suitable for learning flow maps. To achieve this goal, we model the deformation $\phi$ as
$\phi_{t}(x;\theta)=\phi(x,t;\theta)=x+t\mathbf{F}(x,t;I_{f},I_{m},\theta),$ (12)
for all $t\in[-1,1]$, where $\mathbf{F}(x,t;I_{f},I_{m},\theta)$ is a TE-UNet with learnable parameters $\theta$ which also receives a pair of fixed and moving images. The reason behind the specific formulation of Eq. 12 is that we can make sure that at $t=0$ we have $\phi_{0}(x)=x$, hence satisfying the identity initial condition of the ODE (3). In order for $\phi$ to be a valid flow map, it needs to satisfy the semigroup property of Eq. 8. In practice, forcing the model to perfectly satisfy the semigroup property could be cumbersome and expensive, since we need to sample $t$ and $s$ independently and impose Eq. 11 twice, which requires more forward calls from the network. To tackle this issue we propose the following simplified semigroup regularization,
$\mathcal{L}^{t}_{\text{semigroup}}=\|\phi_{2t-1}-\phi_{t}\circ\phi_{t-1}\|_{2}+\|\phi_{2t-1}-\phi_{t-1}\circ\phi_{t}\|_{2},$ (13)
for all $t\in[0,1]$, and we show that it suffices to render $\phi$ a diffeomorphism at all time steps $t\in[-1,1]$. In Eq. 13, .$\|.\|_{2}$ denotes the $L_{2}$-norm, and forces the model to satisfy the semigroup property for $t$ and $s=t-1$. The second term in Eq. 13 is added to impose symmetry and is essential to hold the semi-group property. Similar to all DIR methods we implement the composition operation by trilinear interpolation of grids over each other [20, 10, 43, 17]. From Eq. 13, we can say that a model that minimizes Eq. 13, will satisfy the following rules
$\phi_{2t-1}=\phi_{t}\circ\phi_{t-1},~{}\phi_{2t-1}=\phi_{t-1}\circ\phi_{t}.$ (14)
Proposition 1.
The deformation $\phi(x,t)$ that satisfies the composition rule of Eq. 14 and $\phi(x,0)=x$ is an exponential map, or equivalently, it is a one-parameter diffeomorphism solving an autonomous ODE (3)(see [50]).
Proof.
See Appendix 7.1. ∎
Finally, we can propose the following loss function which is in concordance with Eq. 6:
$\mathcal{L}=\mathbb{E}_{(I_{f},I_{m})\sim\mathcal{D},t\sim Uni(0,1)}\left[\mathcal{L}^{t}_{\text{sim}}+\lambda\mathcal{L}^{t}_{\text{semigroup}}\right],$ (15)
where $Uni(0,1)$ is the uniform distribution on $[0,1]$ and $\mathcal{L}^{t}_{\text{sim}}$ and $\mathcal{L}^{t}_{\text{semigroup}}$ are defined as in Eq. 11 and Eq. 13, respectively. $\lambda$ is the regularization factor controlling the level of diffeomorphism. In the training phase, we randomly sample a pair of images to be registered along with a time step uniformly taken from the interval $[0,1]$ to minimize Eq. 15. The trained model is capable of warping either $I_{f}$ or $I_{m}$ towards the other at any time step. Thus, the model is more versatile compared to the most previous works that provide the deformation in only one direction.
Moreover, note that since the implementation of $\phi_{t}$ as explained above has an inverse for any $t\in[-1,1]$ the network actually learns a continuous group structure of deformations over the interval $[-1,1]$.
It is also worth mentioning that some prior works [31, 32, 30] enforce forward-backward deformation consistency as a regularization term by constructing a time-independent forward deformation only at the trajectory endpoint ($\phi_{1}$), while explicitly modeling the backward deformation ($\phi_{1}^{-1}$). In contrast, SGDIR allows evaluating the deformation field $\phi_{t}$ at arbitrary time steps, without separately modeling the inverse. Consistency between forward and backward deformations, as we prove, is obtained as a natural result of semigroup regularization,
Figure 2: Qualitative comparison of SGDIR with other methods on the OASIS dataset. (visually looks better in color and zoomed in.)
5 Experiments
5.1 Datasets and Metrics
To evaluate SGDIR and compare its performance with other methods, we consider four publicly available datasets on 3D MR brain scans widely used in the literature, namely OASIS [51], CANDI [52], LPBA40 [53], and IXI  [^1] https://brain-development.org/ixi-dataset/. To evaluate the performance of the model and for quantitative comparisons we use Dice score, 95th percentile Hausdorff Distance (HD95), and the percentage of negative Jacobian determinants. We provide the details of datasets, data preparations and evaluation metrics in Appendices 7.2 and 7.4.
5.2 Experimental Settings
In the OASIS dataset, out of 416 subjects, we use 256, 60, 100 subjects for training, validation, and testing, respectively. By random sampling we generate 1000 pairs for training, 100 pairs for validation and 1000 pairs for test. In the CANDI dataset, out of 117 subjects, we use 80, 11, 26 subjects for training, validation, and testing, respectively. We generate 400 pairs for training, 25 pairs for validation, and 300 pairs for test. For the LPBA40 dataset 190 pairs from 20 subjects are used as the training set, 10 pairs from 5 subjects are used as the validation set and 105 pairs from 15 subjects are used for the test set. Also, for the IXI dataset, out of 581 subjects, we use 450, 50, 81 subjects for training, validation, and testing, respectively, from which we create 1500, 100, 300 random pairs for training, validation, and test. We compare the performance of our model with a traditional optimization-based method, SyN [4], as one of the most successful methods for diffeomorphic image registration, learning-based methods including SYMNet [20], VoxelMorph [21], LapIRN [43], DiffuseMorph [27], PULPo [44], TransMorph-diff (TM-diff) [46], and optimization-based methods such as IDIR [8], NODEO [9], NePhi [39], DNVF and its variant Cas-DNVF [10]. We also compare our method with several non-diffeomorphic deformable methods including H-ViT [47], CorrMLP [54], CycleMorph [30], and TransMatch [55]. For SyN, we used the implementation in the DIPY package [56] and used cross-correlation as the loss function following the convention of the VoxelMorph. For other methods, except for DNVF/Cas-DNVF whose implementation was not available at the time of writing this paper, we used the codes available in their corresponding official repositories. The architecture of the UNet along with the time embedding modules is shown in Figure 1. We have provided our full implementation setup in Appendix 7.6.
5.3 Results
Tables 2, 2, 4, 4 provide a comprehensive comparison of the performance of SGDIR with other aforementioned methods on OASIS, CANDI, LPBA40, and IXI, respectively. These tables provide the performance of models in terms of Dice score, HD95 metric, and the percentage of negative Jacobian determinants. All the values are expressed as their average over the test set along with their standard deviation. For DNVF and Cas-DNVF due to unavailability of the implementation the results are directly taken from the corresponding paper. We consider the percentage of negative Jacobian determinants smaller than $10^{-5}\%$ negligible and show it as $\approx 0$ in the tables. Also, Figure 2 illustrates the qualitative comparison between our model and several other models on the OASIS dataset in terms of the warped image, warped segmentations, and the deformation grids. More visual results can be found in Appendix 7.9.
Table 1: Comparison of the performance of SGDIR with other methods on the OASIS dataset. The values inside the parenthesis indicate the standard deviation.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 73.14 (1.32) 3.01 (0.28) 0.0531% (0.0241)
Learning VoxelMorph 75.93 (1.44) 2.31 (0.24) 0.0026% (0.0017)
SYMNet 79.39 (1.74) 1.93 (0.37) 0.0043% (0.0017)
LapIRN 78.80 (1.92) 1.87 (0.23) 0.0011% (0.0003)
DiffuseMorph 76.23 (2.23) 2.34 (0.44) 0.0551% (0.0131)
PULPo 77.36 (1.48) 2.05 (0.32) 0.0137% (0.0085)
CorrMLP 80.78 (1.33) 1.67 (0.33) 0.1162% (0.0310)
H-ViT 81.14 (1.31) 1.53 (0.26) 0.31% (0.012)
TransMatch 80.02 (1.14) 1.87 (0.34) 0.42% (0.093)
CycleMorph 77.32 (2.34) 1.97 (0.33) 0.22% (0.15)
TransMorph-diff 79.84 (1.93) 1.88 (0.23) 0.0012 (0.0006)
Optimization IDIR 72.31 (1.23) 2.75 (0.34) 0.0389% (0.0105)
NODEO 80.86 (1.14) 1.55 (0.12) 0.0105% (0.0041)
NePhi 77.89 (1.34) 1.77 (0.13) 0.0011% (0.0003)
DNVF 79.4 - 0.0015%
Cas-DNVF 81.5 - 0.0036%
Our Method SGDIR 82.23 (1.38) 1.44 (0.16) 0.0009% (0.0007)
Table 2: Comparison of the performance of SGDIR with other methods on the CANDI dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 73.60 (1.15) 2.13 (0.18) 0.0011% (0.0008)
Learning VoxelMorph 75.63 (1.78) 1.78 (0.24) 0.0011% (0.0009)
SYMNet 75.89 (1.13) 1.77 (0.24) 0.0015% (0.0008)
LapIRN 76.23 (1.76) 1.69 (0.21) 0.0004% (0.0002)
DiffuseMorph 75.63 (2.01) 1.83 (0.26) 0.0041% (0.0012)
PULPo 77.97 (1.43) 1.80 (0.17) 0.0097% (0.0011)
CorrMLP 78.88 (1.92) 1.78 (0.19) 0.0018% (0.0007)
H-ViT 80.63 (1.10) 1.66 (0.14) 0.0476% (0.0184)
CycleMorph 77.31 (2.09) 1.55 (0.13) 0.0024% (0.0011)
TransMorph-diff 77.45 (1.39) 1.51 (0.19) 0.0021% (0.0009)
Optimization IDIR 74.42 (1.12) 1.98 (0.11) 0.0113% (0.0081)
NODEO 78.02 (1.52) 1.76 (0.24) $\approx$ 0%
NePhi 78.89 (1.44) 1.73 (0.21) 0.0004% (0.0001)
DNVF 78.6 - 0.0003%
Our Method SGDIR 80.73 (1.25) 1.33 (0.20) $\mathbf{\approx 0}$%
Table 3: Comparison of the performance of SGDIR with other methods on the LPBA40 dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 66.82 (3.17) 5.08 (0.34) 0.0194% (0.0072)
Learning VoxelMorph 70.72 (2.33) 4.02 (0.23) 0.0023% (0.0014)
SYMNet 69.98 (1.37) 4.06 (0.17) 0.0018% (0.0012)
LapIRN 71.88 (1.36) 3.98 (0.26) 0.0011% (0.0005)
DiffuseMorph 68.73 (2.88) 4.13 (0.32) 0.0032% (0.0011)
PULPo 68.34 (1.88) 4.36 (0.45) 0.0052% (0.0014)
CorrMLP 73.12 (1.06) 3.89 (0.21) 0.0315% (0.0021)
H-ViT 74.87 (1.73) 3.73 (0.19) 0.0891% (0.033)
TransMatch 71.52 (1.66) 4.01 (0.27) 0.0981% (0.0113)
CycleMorph 71.03 (1.54) 3.97 (0.32) 0.0413% (0.0087)
TransMorph-diff 72.91 (1.46) 3.83 (0.31) 0.0093% (0.0031)
Optimization IDIR 67.93 (2.03) 4.75 (0.43) 0.0013% (0.0008)
NODEO 72.31 (1.12) 3.96 (0.23) 0.0009% (0.0004)
NePhi 69.39 (1.98) 4.76 (0.74) 0.0103% (0.0045)
Our Method SGDIR 74.23 (2.15) 3.78 (0.13) 0.0007% (0.0003)
Table 4: Comparison of the performance of SGDIR with other methods on the IXI dataset.
Category Model Dice ($\uparrow$) HD95 ($\downarrow$) $|\text{det}(J_{\phi})<0|$ ($\downarrow$)
Traditional SyN 65.14 (2.09) 5.12 (0.66) 0.0201% (0.0096)
Learning VoxelMorph 70.76 (2.32) 4.62 (0.31) 0.0575% (0.0043)
SYMNet 71.21 (1.73) 4.46 (0.27) 0.0084% (0.0033)
LapIRN 72.71 (1.94) 4.24 (0.37) 0.0043% (0.0012)
DiffuseMorph 69.11 (2.14) 4.91 (0.76) 0.0175% (0.0053)
PULPo 69.24 (2.31) 4.83 (0.55) 0.0126% (0.0044)
CorrMLP 75.44 (1.32) 3.76 (0.31) 0.0192% (0.0064)
H-ViT 76.02 (1.45) 3.52 (0.24) 0.23% (0.07)
TransMatch 74.19 (1.24) 3.89 (0.66) 0.87% (0.026)
CycleMorph 70.39 (1.43) 4.73 (0.54) 0.47% (0.013)
TransMorph-diff 72.30 (2.01) 4.08 (0.43) 0.0049% (0.0004)
Optimization IDIR 68.25 (2.28) 5.01 (0.68) 0.0105% (0.0033)
NODEO 76.81 (1.09) 3.11 (0.14) 0.0013% (0.0008)
NePhi 73.81 (1.09) 4.14 (0.38) 0.0093% (0.0027)
Our Method SGDIR 77.98 (1.08) 3.02 (0.12) 0.0012% (0.0009)
Tables 2, 2, 4, 4 indicate that SGDIR outperforms competent diffeomorphic methods both in terms of Dice score, and $|det(J_{\phi})<0|$. SGDIR demonstrates 1.7%, 2.3%, 1.8%, and 1.5% improvements on Dice score with respect to the best evaluated diffeomorphic method over the OASIS, CANDI, LPBA40, and IXI datasets respectively. Among the non-diffeomorphic methods, H-ViT demonstrates the highest performance. The experiments demonstrate the SGDIR achieves competitive results against H-ViT in terms of Dice score while significantly improving the smoothness of deformation grid through $|det(J_{\phi})<0|$. This implies that the imposition of the semigroup property as a regularization in the time continuum of the deformation field can in fact lead to a precise and smooth deformation.
Among the compared methods, DiffuseMorph is also capable of generating a temporal deformation flow. Therefore, we have provided a visual comparison of warped images within 7 time steps for the OASIS dataset in Figure 3. In this example we have used the subject id 1 and 2 of the OASIS dataset as the fixed image and the moving image respectively. More visual results are provided in Appendix 7.9. Also, additional results on another dataset with a different modality is available in Appendix 7.8.
Figure 3: The visual comparison of the continuous deformations of DiffuseMorph and SGDIR for OASIS dataset.
5.4 Ablation Study
Here we investigate the effect of some hyperparameters and design choices along with some consequent properties of SGDIR. More precisely, we investigate the effect of the semigroup regularization factor $\lambda$, the effect of continuous time scale against discrete time, and the topology-preserving property throughout the time interval. Since the behavior of the model is the same for all datasets, we provide the analysis only for the OASIS dataset.
5.4.1 Effect of Regularization Factor
Eq. 13 is responsible for imposing the semigroup property to the model and consequently, make the model produce topology-preserving non-folding deformations. Table 5 provides the effect of changing the factor $\lambda$ for the semigroup regularization on the Dice score and $|det(J_{\phi})<0|$. This table indicates that for a choice of a very large $\lambda$ we could achieve zero-folding deformations with a decrease in the dice score. Decreasing $\lambda$ gives more freedom to the model to increase the similarity of the deformed image and the fixed image in exchange for a larger folding percentage. An interesting property exhibited here is that drastically decreasing $\lambda$ (e.g., $\lambda=1$) is not helpful for achieving the highest Dice score. This indicates that the semigroup property is also helpful in narrowing down the trajectory space of the deformations, helping the model to converge to the more optimal solution.
Table 5: Analysis of the effect of the semigroup regularization factor on the Dice score and the percentage of negative Jacobian determinants over the OASIS dataset.
$\lambda$ Dice ()($\uparrow$)( ↑ ) $|\text{det}(J_{\phi})<0|$ ()($\downarrow$)( ↓ )
$10^{7}$ 76.20 0.0000%
$10^{5}$ 82.11 0.0011%
$10^{4}$ 83.45 3.0691%
$10^{3}$ 82.66 6.0955%
$10^{2}$ 80.01 6.7403%
$10$ 80.23 7.0794%
$1$ 79.80 7.8612%
5.4.2 Diffeomorphism Through Time
Applying the semigroup regularization ensures that the deformation is a diffeomorphism not only at the endpoints, but also at all times $t\in[0,1]$. Figure 4 illustrates how $|det(J_{\phi})<0|$ varies across time. The negative time interval is used for reverse warping; i.e., warping the fixed image towards the moving image. The upper figure illustrates the effect of different values of $\lambda$ on $|det(J_{\phi})<0|$ throughout the time interval, and one can see that a proper factor for the semigroup regularization could lead to an almost perfect diffeomorphism in each time step. The bottom figure, focuses on the performance of the best model. The purple plot and the right y-axis indicate the Dice score of intermediate warped images throughout time and the black plot and the left y-axis indicate the $|det(J_{\phi})<0|$. The shaded areas for both plots show the variance of the corresponding metric. We observe that the semigroup regularization successfully preserves the topology at all time steps, while keeping the Dice score high during the transition.
Figure 4: The percentage of negative Jacobian determinants through time. The upper figure compares the effect of $\lambda$ on the $|det(J_{\phi})<0|$ and the bottom s figure focuses on the best configuration model. The shaded area in the bottom figure indicates the variance across the test set.
5.4.3 Effect of Discrete Time Steps
SGDIR can compute the deformation at an arbitrary time step using a continuous sampling from the time interval $[0,1]$. Here, we investigate the effect of using discrete time steps for training the model. The number of time steps indicates the number of intermediate time steps used during training. For instance, 0 time steps means that only the endpoint are used for registration (at time 1), and 8 time steps means that 8 equidistant intermediate time steps are used during training. $\infty$ is the continuous time sampling as done in our method. At the evaluation phase, we query the model at times 1 and -1 to warp the moving image towards the fixed image, and vice versa. Table 6 demonstrates that increasing the number of time samples in the interval $[0,1]$ improves the performance of the model both in terms of Dice score and $|det(J_{\phi})<0|$, while the best performance is achieved by continuous sampling. Figure 5 in Appendix 7.9 provides an illustrative inspection of this effect as well.
Table 6: Effect of the number time steps on the Dice score and the percentage of negative Jacobian determinants.
# Time Steps Dice $|\text{det}(J_{\phi})<0|$
00 73.12 13.121%
$1$ 73.68 0.152%
$2$ 74.23 0.065%
$4$ 75.03 0.008%
$8$ 75.45 0.002%
$\infty$ 82.23 0.0009%
Table 7: Inference time comparison between SGDIR and other methods.
Model Inference Time (s) ()($\downarrow$)( ↓ )
SYMNet 0.41
VoxelMorph 0.41
TM-diff 0.39
PULPo 0.38
DiffuseMorph 5.77
H-ViT 0.31
NODEO 58.37
IDIR 149.24
NePhi 171.86
SGDIR 0.23
5.5 Computational Analysis
SGDIR is comprised of a rather simple architecture and avoids scaling and squaring integration. Here we have compared the inference time of SGDIR with other methods to evaluate how removing the integration can help the method to have faster inference time. The inference times are calculated for a single pair of brain MR scans. Table 7 compares the inference times. We note that among the learning-based diffeomorphic methods, SGDIR is 1.8 times as fast as SYMNet which employs scaling and squaring during the inference. DiffuseMorph requires us to take the iterative sampling steps of diffusion models and consequently, has much higher inference time. Due to existence of computationally expensive self- and cross-attentions, H-ViT demonstrates inferior inference time comparing to SGDIR. Finally, instance optimization-based methods which need to optimize the deformation grid for separate pairs, demonstrate orders of magnitude higher inference times, which is expected from such methods. More details on the number of parameters of SGDIR and its memory use can be found in Appendix 7.6 and in Table 9.
6 Conclusion
In this paper, we proposed SGDIR, a novel learning-based model for diffeomorphic image registration with a time-embedded UNet architecture and using a semigroup-based regularization. The time-embedded network allows us to compute deformation and consequently warp images at any time point, and the semigroup regularization enables the model to generate diffeomorphic deformations and its inverse. The resulting model is capable of generating smooth warpings from the moving image to the fixed image and vice versa without requiring any integration scheme or using further regularization terms on the smoothness and inverse consistency. We evaluated our model on four benchmark datasets on 3D medical image registration and demonstrated its superiority in various metrics compared to the state-of-the-art methods. In a broader perspective, the proposed framework provides a systematic method for addressing problems which can be modeled as an autonomous ODE.